OpenStudy (anonymous):

write an equation of a line that is perpendicular to the line y= -1/2x-3

OpenStudy (anonymous):

y= -1/2x-3 m1 = -1/2 Since lines are Perpendicular m1 * m2 = -1 m 2 = 2

OpenStudy (anonymous):

and u Need points through which the line Pass Through to Write the Equation..)